# Integral Springer Theorem for Quadratic Lattices under Base Change of Odd Degree

### v1

@@ 1 Introduction
Let $M$ and $N$ be quadratic lattices over $R$. Let $S$ be the integral closure of $R$ in a finite extension $E/F$ of odd degree. Suppose that $S\otimes_{R}N$ can be embedded into $S\otimes_{R}M$ over $S$. Can we embed $N$ into $M$ over $R$
@@ 2.1 Embedding criterion and integral Springer theorem in the local case
Question 1.1 has a positive answer for the complete discrete valuation ring $R$ (where $2$ is invertible). That is, for any quadratic lattices $M$ and $N$ over $R$, $N$ can be embedded into $M$ over $R$ if and only if $S\otimes_{R}N$ can be embedded into $S\otimes_{R}N$ over $S$, where $S$ is the integral closure of $R$ in a finite extension $E/F$ of odd degree.
@@ 2.2 Uniqueness of Jordan splittings and cancellation law
for $1\leq i\leq t$. Comparing the dimensions, we obtain $F\otimes_{R}L_{1}\perp\cdots\perp F\otimes_{R}L_{i}\cong F\otimes_{R}K_{1}\perp\cdots\perp F\otimes_{R}K_{i}$ for each $i$. By the Witt cancellation theorem for quadratic spaces ([28, 42:16. Theorem]), we have $F\otimes_{R}L_{i}\cong F\otimes_{R}K_{i}$ for each $i$. Hence $L_{i}\cong K_{i}$ for each $i$, by Corollary 2.3. ∎
@@ 3.2 Sets of transporters and their spinor norms
Since the number of Jordan components of $M^{\prime}$ is $m-1$, the result now follows by induction.
@@
$M^{\prime\prime}=\begin{cases}\pi^{(s_{3}-s_{1})/2}M_{1}\perp\pi^{(s_{3}-s_{2}-1)/2}M_{2}\perp M_{3}\perp\cdots\perp M_{m}\ &\text{if $s_{1}\equiv s_{3}\pmod{2}$},\\ \pi^{(s_{3}-s_{1}-1)/2}M_{1}\perp\pi^{(s_{3}-s_{2})/2}M_{2}\perp M_{3}\perp\cdots\perp M_{m}\ &\text{if $s_{1}\equiv s_{3}+1\pmod{2}$}.\end{cases}$
@@
$M^{\prime\prime}:=\begin{cases}\pi^{(s_{3}-s_{1})/2}M_{1}\perp\pi^{(s_{3}-s_{2}-1)/2}M_{2}\perp M_{3}\perp\cdots\perp M_{m}\ \ \ &\text{if $s_{1}\equiv s_{3}\pmod{2}$},\\ \pi^{(s_{3}-s_{1}-1)/2}M_{1}\perp\pi^{(s_{3}-s_{2})/2}M_{2}\perp M_{3}\perp\cdots\perp M_{m}\ \ \ &\text{if $s_{2}\equiv s_{3}\pmod{2}$}.\end{cases}$
@@ 3.3 Norm principles for integral spinor norms
$\mathfrak{M}_{i}:=\{Q(x_{i})\,|\,x_{i}\text{ is a norm generator of }M_{i}\}\quad\text{ for each }1\leq i\leq t$
@@
is a Jordan splitting of $\widetilde{M}$ with $x_{i}\in\widetilde{M_{i}}$ and $Q(x_{i})S=\mathfrak{s}(\widetilde{M_{i}})$ for $1\leq i\leq t$. Since
@@
$\theta(X(M/N))=\theta(X(M^{\prime}/N^{\prime}))\cdot\theta(\mathrm{O}(M))\ \ \ \text{and}\ \ \ \theta(X(\widetilde{M}/\widetilde{N}))=\theta(X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}}))\cdot\theta(\mathrm{O}(\widetilde{M})).$
@@
$\widetilde{N}\subseteq\pi\widetilde{M_{1}}\perp\cdots\perp\widetilde{M_{t}}=\widetilde{M^{*}}\ \ \ \text{and}\ \ \ X(\widetilde{M}/\widetilde{N})=X(\widetilde{M^{*}}/\widetilde{N}).$
@@
Since $2$ is invertible in $R$, the residue degree $[k_{E}:k_{F}]$ is odd. Thus, the ramification index $e$ of $E/F$ is even if $[E:F]=e[k_{E}:k_{F}]$ is even.
@@
$\begin{split}&X(\widetilde{M}/\widetilde{N})=X(\pi_{E}^{e/2}\widetilde{M_{1}}\perp(\widetilde{N_{1}}\perp\widetilde{B_{2}})\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{1}}\perp\cdots\perp\widetilde{N_{n}})\\ =\;&X(\pi_{E}^{e/2}\widetilde{M_{1}}\perp\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\cdot\mathrm{O}(\pi_{E}^{e/2}\widetilde{M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}}).\end{split}$
@@
$\begin{split}&X(\widetilde{M}/\widetilde{N})=X(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp(\widetilde{N_{1}}\perp\widetilde{B_{2}})\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{1}}\perp\cdots\perp\widetilde{N_{n}})\\ =\;&X(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{B_{2}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\cdot\mathrm{O}(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}})\end{split}$ (3.3.1)
@@
Suppose that $E\otimes_{S}(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{B_{2}})$ is isotropic over $E$. Since $F\otimes_{R}(M_{1}\perp B_{2})$ is anisotropic (by Corollary 3.3), $[E:F]$ must be even by the classical Springer theorem, and it follows that $e$ is even. Then $\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}$ and $\widetilde{B_{2}}$ have the same scale,
@@
Now assume that $E\otimes_{S}(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{B_{2}})$ is anisotropic over $E$. Since $t_{2}\geq t_{1}+1=s_{1}+2$, one obtains that
$\begin{split}&X(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\\ =\;&X(\pi_{E}^{e}\widetilde{M_{1}}\perp\pi_{E}^{[e/2]}\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\end{split}$ (3.3.2)
@@
$\begin{split}&X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}})=X(\widetilde{\pi M_{1}}\perp\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\\ =\;&X(\pi_{E}^{e}\widetilde{M_{1}}\perp\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}})\\ =\;&X(\pi_{E}^{e}\widetilde{M_{1}}\perp\pi_{E}^{[e/2]}\widetilde{B_{2}}\perp\widetilde{M_{3}}\perp\cdots\perp\widetilde{M_{m}}/\widetilde{N_{2}}\perp\cdots\perp\widetilde{N_{n}}).\end{split}$ (3.3.3)
@@
$N_{E/F}\big{(}\theta(\mathrm{O}(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}}))\big{)}\subseteq N_{E/F}\big{(}\theta(\mathrm{O}(\pi_{E}^{e}\widetilde{M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}}))).$ (3.3.4)
@@
$\begin{split}&N_{E/F}\big{(}\theta(X(\widetilde{M}/\widetilde{N}))\big{)}=N_{E/F}\big{(}\theta(X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}}))\cdot\theta(\mathrm{O}(\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}}))\big{)}\\ \subseteq\;&N_{E/F}\big{(}\theta(X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}}))\cdot\theta(\mathrm{O}(\widetilde{\pi M_{1}}\perp\widetilde{M_{2}}\perp\cdots\perp\widetilde{M_{m}}))\big{)}=N_{E/F}\big{(}\theta(X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}}))\cdot\theta(\mathrm{O}(\widetilde{M^{*}}))\big{)}.\end{split}$
@@
Now assume $e$ is even. Then $\pi_{E}^{[(e+1)/2]}\widetilde{M_{1}}$ and $\widetilde{M_{2}}$ have the same scale, which is strictly contained in $\mathfrak{s}(\widetilde{M_{3}})$. Thus, by Proposition 3.8, (3.3.4) follows from the following relation:
@@
To finish the proof of (3.3.5), it suffices to show that $N_{E/F}(S^{\times})\subseteq(F^{\times})^{2}$. For any $x\in R$ (resp. $x\in S$), let $\bar{x}\in k_{F}$ (resp. $\bar{x}\in k_{E}$) denote its canonical image in the residue field of $R$ (resp. $S$). For any $\alpha\in S^{\times}$, there exists an element $\beta\in R^{\times}$ such that $\bar{\alpha}^{[k_{E}:k_{F}]}=\bar{\beta}$, because $k_{E}/k_{F}$ is purely inseparable. Since $[k_{E}:k_{F}]$ is odd, $\bar{\alpha}\bar{\beta}^{-1}$ is a square in $k_{E}$. Hence, by Hensel’s lemma, $\alpha=\beta\gamma^{2}$ for some $\gamma\in S^{\times}$. Thus
@@
If $\mathop{{{\perp}}}\limits_{s_{j}<t_{1}}(E\otimes_{S}\widetilde{M_{j}})$ is isotropic over $E$, then
@@
If $\mathop{{{\perp}}}\limits_{s_{j}<t_{1}}(E\otimes_{S}\widetilde{M_{j}})$ is anisotropic over $E$, then by Proposition 3.7,
$X(\widetilde{M}/\widetilde{N})=X\big{(}P_{1}\perp P_{2}\perp(\mathop{\perp}\limits_{s_{j}>t_{1}}\widetilde{M_{j}})/\widetilde{N}\big{)}$
@@
$X\big{(}P_{1}\perp P_{2}\perp(\mathop{\perp}\limits_{s_{j}>t_{1}}\widetilde{M_{j}})/\widetilde{N})\subseteq X(\widetilde{M^{{\ddagger}}}/\widetilde{N}\big{)}\subseteq X(\widetilde{M}/\widetilde{N})$
by definition, one concludes $X(\widetilde{M^{{\ddagger}}}/\widetilde{N})=X(\widetilde{M}/\widetilde{N})$. The result follows from induction on $\mathfrak{s}(M)$. ∎
@@ 4.1 Adelic language and genus theory
$\mathrm{O}(V)\big{\backslash}\mathrm{O}_{\mathbf{A}}(V)\big{/}\mathrm{O}_{\mathbf{A}}(M)\overset{\sim}{\longrightarrow}\{\mathrm{cls}(L)\,|\,L\in\mathrm{gen}(M)\}$
@@
$\begin{split}\mathrm{O}^{+}(V)\big{\backslash}\mathrm{O}^{+}_{\mathbf{A}}(V)\big{/}\mathrm{O}_{\mathbf{A}}^{+}(M)&\overset{\sim}{\longrightarrow}\{\mathrm{cls}^{+}(L)\,|\,L\in\mathrm{gen}(M)\}\,,\\ \mathrm{O}^{+}(V)\underline{\sigma}\mathrm{O}^{+}_{\mathbf{A}}(M)&\longmapsto\mathrm{cls}^{+}(\underline{\sigma}M).\end{split}$
@@
$\begin{split}\big{(}\mathrm{O}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V)\big{)}\big{\backslash}\mathrm{O}_{\mathbf{A}}(V)\big{/}\mathrm{O}_{\mathbf{A}}(M)&\overset{\sim}{\longrightarrow}\{\mathrm{spn}(L)\,|\,L\in\mathrm{gen}(M)\}\,,\\ \big{(}\mathrm{O}^{+}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V)\big{)}\big{\backslash}\mathrm{O}^{+}_{\mathbf{A}}(V)\big{/}\mathrm{O}^{+}_{\mathbf{A}}(M)&\overset{\sim}{\longrightarrow}\{\mathrm{spn}^{+}(L)\,|\,L\in\mathrm{gen}(M)\}.\end{split}$ (4.1.1)
@@ 4.2 Spinor genera and classes containing a given sublattice
$\begin{split}\mathrm{O}_{\mathbf{A}}(M)&\subseteq X_{\mathbf{A}}(M/N)=X_{\mathbf{A}}(M/N)\mathrm{O}_{\mathbf{A}}(M)\,,\\ \text{and }\ \mathrm{O}^{+}_{\mathbf{A}}(M)&\subseteq X^{+}_{\mathbf{A}}(M/N)=X^{+}_{\mathbf{A}}(M/N)\mathrm{O}^{+}_{\mathbf{A}}(M).\end{split}$ (4.2.1)
@@
$\begin{split}\{\mathrm{spn}(L)\,|\,L\in\mathrm{gen}(M),\;N\hookrightarrow\mathrm{spn}(L)\}&\cong(\mathrm{O}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V))\big{\backslash}\mathrm{O}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V)X_{\mathbf{A}}(M/N)\big{/}\mathrm{O}_{\mathbf{A}}(M),\\ \{\mathrm{spn}^{+}(L)\,|\,L\in\mathrm{gen}(M),\;N\hookrightarrow\mathrm{spn}^{+}(L)\}&\cong(\mathrm{O}^{+}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V))\big{\backslash}\mathrm{O}^{+}(V)\mathrm{O}^{\prime}_{\mathbf{A}}(V)X^{+}_{\mathbf{A}}(M/N)\big{/}\mathrm{O}^{+}_{\mathbf{A}}(M).\end{split}$
@@
[…]
@@ 5.1 Norm principles for global integral and adelic spinor norms
Then it would imply $[A:A^{+}]=[\theta(\mathrm{O}(M)):\theta(\mathrm{O}^{+}(M))]\leq 2$. But this is not the case, because $p_{1},p_{2}$ and their quotient $p_{1}/p_{2}$ all lie in $A\backslash A^{+}$. Therefore (5.1.1) fails in this case.
@@
$N_{E/F}\big{(}\theta(\mathrm{O}^{+}(\widetilde{L}))\big{)}\supseteq\{\alpha\in R^{\times}:\alpha>0\ \text{in $F_{v}$ for all real places $v$ of $F$}\}.$
Therefore the fundamental unit $2-\sqrt{3}$ of $R$ belongs to $N_{E/F}\big{(}\theta(\mathrm{O}^{+}(\widetilde{L}))\big{)}$. Therefore,
@@
We only give a proof of the first inclusion, since the other can be proved similarly. Consider an arbitrary $\underline{\sigma}=(\sigma_{\mathfrak{q}})\in X_{\mathbf{A}}(\widetilde{M}/\widetilde{N})$. By definition, we have
@@
If $E/F$ is a purely inseparable extension of prime degree $l\neq 2$, then $E=F(\sqrt[l]{b})$ for some $b\in F^{\times}\setminus(F^{\times})^{l}$, and hence $F_{\mathfrak{p}}\otimes_{F}E\cong F_{\mathfrak{p}}[x]/(x^{l}-b).$ Since the polynomial $x^{l}-b$ either stays irreducible or factors as $(x-c)^{l}$ for some $c\in F_{\mathfrak{p}}$ over $F_{\mathfrak{p}}$, the Jacobson radical of $F_{\mathfrak{p}}\otimes_{F}E$ is either trivial or generated by the canonical image of $x-c$. It then follows from [4, Chapter VI, § 8.2, Proposition 2 (b)] that there is a unique $\mathfrak{q}\in\mathrm{Spm}(S)$ lying above $\mathfrak{p}$ and that
$[E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=[E_{\mathfrak{q}}:F_{\mathfrak{p}}]=\begin{cases}l\quad&\text{ if }x^{l}-b\text{ is irreducible over }F_{\mathfrak{p}}\,,\\ 1\quad&\text{ otherwise}\,.\end{cases}$
@@
are clearly compatible if $[E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=l$. Since $l$ is odd, these two norms are compatible modulo squares if $[E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=1$.

### v2

@@ 1 Introduction
Let $M$ and $N$ be quadratic lattices over $R$. Let $S$ be the integral closure of $R$ in a finite extension $E/F$ of odd degree. Suppose that $S\otimes_{R}N$ can be embedded into $S\otimes_{R}M$ over $S$. Can we embed $N$ into $M$ over $R$ ?
@@ 2.1 Embedding criterion and integral Springer theorem in the local case
Question 1.1 has a positive answer for the complete discrete valuation ring $R$ (where $2$ is invertible). That is, for any quadratic lattices $M$ and $N$ over $R$, NN can be embedded into $M$ over $R$ if and only if $S\otimes_{R}N$ can be embedded into $S\otimes_{R}N$ over $S$, where $S$ is the integral closure of $R$ in a finite extension $E/F$ of odd degree.
@@ 2.2 Uniqueness of Jordan splittings and cancellation law
for $1\leq i\leq t$. Comparing the dimensions, we obtain $F\otimes_{R}L_{1}\perp\cdots\perp F\otimes_{R}L_{i}\cong F\otimes_{R}K_{1}\perp\cdots\perp F\otimes_{R}K_{i}$ for each ii. By the Witt cancellation theorem for quadratic spaces ([28, 42:16. Theorem]), we have $F\otimes_{R}L_{i}\cong F\otimes_{R}K_{i}$ for each ii. Hence $L_{i}\cong K_{i}$ for each $i$, by Corollary 2.3. ∎
@@ 3.2 Sets of transporters and their spinor norms
Since the number of Jordan components of $M^{\prime}$ is m1m-1, the result now follows by induction.
@@
[…]
@@
[…]
@@ 3.3 Norm principles for integral spinor norms
[…]
@@
[…]
@@
$).\theta(X(M/N))=\theta(X(M^{\prime}/N^{\prime}))\cdot\theta(\mathrm{O}(M))\ \ \ \text{and}\ \ \ \theta(X(\widetilde{M}/\widetilde{N}))=\theta(X(\widetilde{M^{\prime}}/\widetilde{N^{\prime}}))\cdot\theta(\mathrm{O}(\widetilde{M}$)).
@@
[…]
@@
Since $2$ is invertible in $R$, the residue degree $k_{E}:k_{F}]$ is odd. Thus, the ramification index $e$ of $E/F$ is even if $E:F]=e[k_{E}:k_{F}]$ is even.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
To finish the proof of (3.3.5), it suffices to show that $N_{E/F}(S^{\times})\subseteq(F^{\times})^{2}$. For any $x\in R$ (resp. $x\in S$), let $\bar{x}\in k_{F}$ (resp. $\bar{x}\in k_{E}$) denote its canonical image in the residue field of $R$ (resp. $S$). For any $\alpha\in S^{\times}$, there exists an element $\beta\in R^{\times}$ such that $\bar{\alpha}^{[k_{E}:k_{F}]}=\bar{\beta}$, because $k_{E}/k_{F}$ is purely inseparable. Since $k_{E}:k_{F}]$ is odd, $\bar{\alpha}\bar{\beta}^{-1}$ is a square in $k_{E}$. Hence, by Hensel’s lemma, $\alpha=\beta\gamma^{2}$ for some $\gamma\in S^{\times}$. Thus
@@
[…]
@@
[…]
@@
[…]
@@ 4.1 Adelic language and genus theory
[…]
@@
[…]
@@
[…]
@@ 4.2 Spinor genera and classes containing a given sublattice
[…]
@@
[…]
@@
${G(F)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${\mathrm{O}^{+}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\theta}$${F^{*}/F^{*2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G_{\mathbf{A}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\mathbf{A}}}$${\mathrm{O}_{\mathbf{A}}^{+}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\theta_{\mathbf{A}}}$${\mathbf{I}_{F}/\mathbf{I}_{F}^{2}}$
@@ 5.1 Norm principles for global integral and adelic spinor norms
[…]
@@
[…]
@@
[…]
@@
If $E/F$ is a purely inseparable extension of prime degree $l\neq 2$, then $E=F(\sqrt[l]{b})$ for some $b\in F^{\times}\setminus(F^{\times})^{l}$, and hence $F_{\mathfrak{p}}\otimes_{F}E\cong F_{\mathfrak{p}}[x]/(x^{l}-b).$ Since the polynomial $x^{l}-b$ either stays irreducible or factors as $(x-c)^{l}$ for some $c\in F_{\mathfrak{p}}$ over $F_{\mathfrak{p}}$, the Jacobson radical of $F_{\mathfrak{p}}\otimes_{F}E$ is either trivial or generated by the canonical image of xcx-c. It then follows from [4, Chapter VI, § 8.2, Proposition 2 (b)] that there is a unique $\mathfrak{q}\in\mathrm{Spm}(S)$ lying above $\mathfrak{p}$ and that
$E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=[E_{\mathfrak{q}}:F_{\mathfrak{p}}]=\begin{cases}l\quad&\text{ if }x^{l}-b\text{ is irreducible over }F_{\mathfrak{p}}\,,\\ 1\quad&\text{ otherwise}\,.\end{cases}$
@@
are clearly compatible if $E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=l$. Since $l$ is odd, these two norms are compatible modulo squares if $E_{(\mathfrak{p})}:F_{\mathfrak{p}}]=1$.